# Optimizing Treatment Allocation in the Presence of Interference

### v1

@@ 3 Problem setup
Overlap and strong ignorability. Moreover, we introduce the overlap and strong ignorability assumptions in the presence of interference (Forastiere et al., 2021; Jiang & Sun, 2022): $\exists\ \delta\in(0,1)$ such that $\delta<p(T_{i}=1|\mathbf{X}_{i}=\mathbf{x}_{i},\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}})<1-\delta$ and $Y_{i}(T_{i}=t,Z_{i}=z)\perp\!\!\!\perp$ $T_{i},Z_{i}$ $|\mathbf{X}_{i},\mathbf{X}_{\mathcal{N}_{i}}$, $\forall t\in\mathcal{T}$, $z\in\mathcal{Z}$, $\mathbf{X}_{i}\in\mathcal{X}$ and $\mathbf{X}_{\mathcal{N}_{i}}=\{\mathbf{X}_{j}\}_{j\in\mathcal{N}_{i}}$. However, the traditional ignorability assumption no longer holds: $Y_{i}(T_{i}=t)\centernot{\perp\!\!\!\perp}$ $T_{i}$ $|\mathbf{X}_{i}$ $\forall t\in\mathcal{T}$ and $\mathbf{X}_{i}\in\mathcal{X}$.
Treatment effect definitions.  Based on the introduced assumptions and notation, the marginal individual treatment effect (MITE) is defined as:
$\tau_{i}(z)=\mathbb{E}\bigl{[}Y_{i}(1,z)-Y_{i}(0,z)|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\bigr{]},$ (1)
with $Y_{i}(t,z)=Y_{i}(T_{i}=t,Z_{i}=z)$. Equation 1 shows that MITE is dependent on the value of $Z_{i}$. Note that the MITE differs from the traditional individual treatment effect (ITE) as the latter does not take into account graph information and is traditionally defined as follows:
$\tau_{i,trad}=\mathbb{E}\bigl{[}Y_{i}(T_{i}=1)-Y_{i}(T_{i}=0)|\mathbf{x}_{i}\bigr{]}.$ (2)
$\tau_{i,trad}=\mathbb{E}_{(\mathbf{X}_{\mathcal{N}_{i}},Z_{i})}\bigl{[}\mathbb{E}\left[Y_{i}(T_{i}=1,Z_{i}=z)-Y_{i}(T_{i}=0,Z_{i}=z)|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\right]\bigr{]}.$ (3)
In this definition, we marginalize out $\mathbf{X}_{\mathcal{N}_{i}}$ and $Z_{i}$. This is what classical causal inference techniques would estimate in the presence of interference.
@@
$\delta_{i}(t,z)=\mathbb{E}\bigl{[}Y_{i}(t,z)-Y_{i}(t,0)|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\bigr{]}.$ (4)
@@
$\omega_{i}(t,z)=\mathbb{E}\bigl{[}Y_{i}(t,z)-Y_{i}(0,0)|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\bigr{]}.$ (5)
@@
$\mathbf{t^{*}}=\underset{\mathbf{t}}{\arg\max}\sum_{i\in\mathcal{N}}\omega_{i}(t_{i},z_{i})\text{ s.t. }\sum_{i\in\mathcal{N}}t_{i}\leq k.$ (7)
Both $t_{i}$ and $z_{i}$ are decision variables and can be derived from $\mathbf{t}$. Now, if we have a model that can estimate the ITTE for a given $\mathbf{t}$, we can use (the sum of) these estimates as our objective function to find an optimal treatment allocation $\mathbf{t^{*}}$.
@@ 4 OTAPI
In UM, the focus lies on finding the entities with the highest expected ITE. The goal in UM is essentially the same as (7). To maximize TTE, entities are ranked according to estimated ITE, and the top $k$ entities are given treatment. Alternatively, integrated approaches exist that rank entities directly without estimating ITE first (Devriendt et al., 2020). Nevertheless, these approaches are not tailored to scenarios involving network interference because they rely solely on individual information. When network interference plays a significant role, ITE estimates may be biased, potentially resulting in suboptimal treatment allocation, where entities with substantial network effects might be overlooked and connected entities might all be treated while treating just one of them might have a similar effect in terms of TTE (Forastiere et al., 2021). Moreover, each treatment allocation vector $\mathbf{t}$ may result in a different ITTE $\omega_{i}$ in the presence of interference. Therefore, simply ranking entities according to ITTE no longer works because an entity’s ITTE depends on the set of other entities that receive treatment.
Classic IM methods focus on modeling network effects, often using diffusion processes to simulate interference dynamics. These require computationally expensive MC simulations to estimate TTE. These TTE estimates lack any form of individualized treatment effect estimation and are commonly based on assumptions regarding the applicable diffusion process. By leveraging heuristics, the TTE is maximized without having to run simulations for every possible treatment allocation.
OTAPI addresses these limitations by combining the strengths of these two methods. It consists of two components: (i) a relational causal estimator that considers both individualized susceptibility to treatment and interference to estimate ITTE (Section 4.1), and (ii) an allocation algorithm that utilizes the predictions made by the trained causal model to allocate treatments optimally (Section 4.2).
@@ 4.1 Estimating treatment effects in the presence of interference
In order to optimally allocate treatment, accurate estimates of the ITTE $\omega_{i}(t_{i},z_{i})$ are required. To estimate unbiased treatment effects, we employ causal machine learning. OTAPI is agnostic to the estimator employed, but its performance directly depends on the accuracy of the estimates.
Given the assumptions in Section 3, we utilize a causal estimator $\mathcal{M}(\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},t_{i},z_{i})$ which provides unbiased estimates for the potential outcomes $\mathbb{E}\bigl{(}Y_{i}(T_{i}=t_{i},Z_{i}=z_{i})|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\bigr{)}$. Since an entity can have an arbitrary number of neighbors, a GNN is employed to learn an aggregation of the neighbors’ features. Jiang & Sun (2022) shows that if the distributions $p(T_{i}|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}})$, the classic propensity score, and $p(Z_{i}|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},t_{i})$, the peer exposure score, are biased, the estimates for the treatment effects will also be biased. Given that treatments are typically not allocated randomly in observational data due to confounding, controlling for these biased distributions is necessary.
To address the challenge of confounding bias in the presence of interference, representation balancing techniques can be applied (Ma & Tresp, 2021; Jiang & Sun, 2022; Shalit et al., 2017). This leads to more accurate estimates of the ITTE. In the experiments presented in Section 5, we adopt NetEst (Jiang & Sun, 2022), a GNN-based causal estimator that employs an adversarial approach to balance representations. Nevertheless, alternative relational causal estimators could also be employed.
@@ 4.2 Allocating treatments optimally
Assuming we can compute unbiased ITTE estimates for all entities as $\hat{\omega}_{i}(t_{i},z_{i})=\mathcal{M}(\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},t_{i},z_{i})$- $\mathcal{M}(\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},0,0)$, we can plug these estimates into the objective (Equation 7). As already explained, the strategy of ranking entities according to estimated ITTE $\hat{\omega}_{i}$ no longer works. To address this, we rely on heuristics developed in the IM literature and plug data-driven estimates of TTE into these algorithms. Generally, these heuristics maximize TTE by iteratively improving a (set of) solution(s). Some examples include the greedy algorithm (Kempe et al., 2003), the genetic algorithm (Bucur & Iacca, 2016), and simulated annealing (Jiang et al., 2011). OTAPI can use any of these algorithms to optimize treatment allocation.
@@ 5.1.1 Data
Synthetic data.  For each synthetic dataset, 3 separate networks (train, validation, and test) are generated according to the Barabási-Albert random network model (Barabási & Albert, 1999). Each network consists of 5,000 nodes. The hyperparameter which determines the number of edges attached from each new node is set to $m=2$. The data-generating process (DGP) is inspired by Jiang & Sun (2022).
For each node, a feature vector of $10$ features is randomly sampled according to a standard normal distribution:
$\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$
with $\mathbf{I}$ the identity covariance matrix. We first generate the parameters that will be used later on in the DGP:
@@
Where Unif is the uniform distribution. These parameters are used to quantify the effect of $\mathbf{X_{i}}$ on ${T_{i}}$, $\mathbf{X_{i}}$ on $Y_{i}$, and the effect of $\mathbf{T_{i}}$ on $Y_{i}$ respectively. The treatment $t_{i}$ is generated as follows:
$&t_{i}\sim\text{Bernoulli}(p_{t_{i}});&p_{t_{i}}=\beta_{XT}\cdot p_{i}+\beta_{{\mathcal{N}}T}\cdot p_{\mathcal{N}_{i}},\\ \end{aligned}$
@@
with $p_{i}=\sigma(\mathbf{w}^{XT}\cdot\mathbf{x}_{i})$ and $p_{\mathcal{N}_{i}}=\frac{\sum_{j\in\mathcal{N}_{i}}p_{j}}{|\mathcal{N}_{i}|}$. Here, $\sigma$ is the sigmoid function. Note that confounding is introduced by generating propensity scores $p_{i}$ and $p_{\mathcal{N}_{i}}$ based on $\mathbf{x}_{i}$ and $\mathbf{x}_{\mathcal{N}_{i}}$, respectively. Now, we describe how the outcome $y_{i}$ is generated:
[…]
@@
where $h_{i}=\sigma(\mathbf{w}^{TY}\cdot\mathbf{x}_{i})+b_{TY}$, $u_{i}=\sigma(\mathbf{w}^{XY}\cdot\mathbf{x}_{i})$, and $u_{\mathcal{N}_{i}}=\frac{\sum_{j\in\mathcal{N}_{i}}u_{j}}{|\mathcal{N}_{i}|}$. Recall that $z_{i}$ is the ratio of $i$’s neighbors that received treatment $z_{i}=\frac{\sum_{j\in\mathcal{N}_{i}}t_{j}}{|\mathcal{N}_{i}|}$. $h_{i}$ is used to make the treatment effect heterogeneous, i.e., dependent on the features. Note that the individual effect is 100% correlated with the spillover effect. This choice is motivated by the fact that if someone is more/less susceptible to treatment, they will most likely also be more/less susceptible to the treatment of neighbors. The following setting for the parameters was used for all experiments:
[…]
The intuition behind these parameters is that the individual features and treatments will always be more impactful than those of neighbors. $\beta_{spillover}$ is varied between 0 and 0.7 across experiments.
Semi-synthetic data.   Following Jiang & Sun (2022) and Guo et al. (2020), we use two real social network datasets: BlogCatalog (BC) and Flickr. We follow the same procedure as Jiang & Sun (2022) to generate the networks and features from the Flickr and BC datasets. Some summary statistics of these datasets are given in Table 1.
Flickr BC
Train Validation Test Train Validation Test
Nodes 2,482 2,461 2,358 1,716 1,696 1,784
Edges 46,268 14,419 23,529 17,937 25,408 14,702
Features 12,047 12,047 12,047 8,189 8,189 8,189
Table 1: Flickr and BC datasets
These networks are split using METIS (Karypis & Kumar, 1998) which results in three separate networks: a train, validation, and test set. We directly use the same split from Jiang & Sun (2022). Given the large number of (sparse) features, we leverage LDA (Blei et al., 2003) to reduce this number to 10 for both datasets. Now, the same DGP as in the synthetic datasets is used. For the semi-synthetic datasets, $\beta_{spillover}=0.3$. All other parameters are the same.
@@
Figure 3: NetEst architecture

### v2

@@ 3 Problem setup
Moreover, we adopt the overlap and strong ignorability assumptions in the presence of interference (Forastiere et al., 2021; Jiang and Sun, 2022):
Overlap and strong ignorability. $\exists\ \delta\in(0,1)$ such that $\delta<p(T_{i}=1|\mathbf{X}_{i}=\mathbf{x}_{i},\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}})<1-\delta$ and $Y_{i}(T_{i}=t,Z_{i}=z)\perp\!\!\!\perp T_{i},Z_{i}\mid\mathbf{X}_{i},\mathbf{X}_{\mathcal{N}_{i}},\forall t\in\mathcal{T},z\in\mathcal{Z},\mathbf{X}_{i}\in\mathcal{X}$ and $\mathbf{X}_{\mathcal{N}_{i}}=\{\mathbf{X}_{j}\}_{j\in\mathcal{N}_{i}}$. However, the traditional ignorability assumption no longer holds: $Y_{i}(T_{i}=t)\centernot{\perp\!\!\!\perp}$ $T_{i}\mid\mathbf{X}_{i}$ $\forall t\in\mathcal{T}$ and $\mathbf{X}_{i}\in\mathcal{X}$. Under the stated assumptions, the causal estimands defined in the next paragraph are identifiable (Forastiere et al., 2021), meaning that they can be expressed as functions of an observed data distribution. Consequently, these causal effects can be estimated from data. A proof of identifiability is provided in Appendix A.
Treatment effect definitions.We now define our causal estimands of interest. The Marginal Individual Treatment Effect (MITE) is defined as:
$\tau_{i}(z)=\mathbb{E}\bigl{[}Y_{i}(1,z)-Y_{i}(0,z)\mid\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\bigr{]}.$ (1)
Equation 1 shows that MITE is dependent on the value of $Z_{i}=z$. The difference between MITE and the traditional ITE is clarified in Appendix B.
@@
$\delta_{i}(t,z)=\mathbb{E}\bigl{[}Y_{i}(t,z)-Y_{i}(t,0)\mid\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\bigr{]}.$ (2)
@@
$\omega_{i}(t,z)=\mathbb{E}\bigl{[}Y_{i}(t,z)-Y_{i}(0,0)\mid\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\bigr{]}.$ (3)
@@
$\mathbf{t^{*}}=\underset{\mathbf{t}}{\arg\max}\sum_{i\in\mathcal{N}}\omega_{i}(t_{i},z_{i})\quad\text{s.t.}\quad\sum_{i\in\mathcal{N}}t_{i}\leq k.$ (5)
Both $t_{i}$ and $z_{i}$ are decision variables and can be derived from $\mathbf{t}$. Now, if we have a model that can estimate the ITTEs for a given $\mathbf{t}$, we can use (the sum of) these estimates as our objective function to find an optimal treatment allocation $\mathbf{t^{*}}$.
@@ 4 OTAPI
In UM, the focus lies on finding the entities with the highest expected ITE. The goal in UM is essentially the same as in Equation 5. To maximize TTE, entities are ranked according to estimated ITE, and the top $k$ entities are given treatment. Alternatively, integrated approaches exist that rank entities directly without estimating ITE first (Devriendt et al., 2020). Nevertheless, these approaches are not tailored to scenarios involving network interference because they rely solely on individual information. When network interference plays a significant role, ITE estimates may be biased, potentially resulting in suboptimal treatment allocation, where entities with substantial network effects might be overlooked and connected entities might all be treated while treating just one of them might have a similar effect in terms of TTE (Forastiere et al., 2021). Moreover, each treatment allocation vector may result in a different ITTE in the presence of interference. Therefore, simply ranking entities according to ITTE no longer works because an entity’s ITTE depends on the set of other entities that receive treatment.
Classic IM methods focus on modeling network effects, often using diffusion processes to simulate interference dynamics. These require computationally expensive MC simulations to estimate TTE, also defined as influence spread in the IM literature. Because it is often infeasible to estimate the TTE for every possible combination of treatment allocations in larger graphs, heuristics are used to limit the number of treatment allocation combinations for which TTE is estimated (Kempe et al., 2003; Bucur and Iacca, 2016). These TTE estimates lack any form of individualized treatment effect estimation and are commonly based on assumptions regarding the applicable diffusion process.
OTAPI addresses these limitations by combining the strengths of these two methods. It consists of two components: (i) a relational causal estimator that considers both individualized susceptibility to treatment and interference to estimate ITTE (Section 4.1), and (ii) an allocation algorithm that uses the predictions made by the trained causal model to allocate treatments optimally (Section 4.2). A high-level pseudocode description of OTAPI is provided in Appendix D.
@@ 4.1 Estimating treatment effects in the presence of interference
In order to optimally allocate treatment, accurate estimates of the ITTE are required. To estimate unbiased treatment effects, we employ causal machine learning. OTAPI is agnostic to the estimator employed, but its performance directly depends on the accuracy of the estimates.
Given the assumptions in Section 3, we use a causal estimator $\mathcal{M}(\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},t_{i},z_{i})$ which provides unbiased estimates for the potential outcomes $\mathbb{E}\left[Y_{i}(T_{i}=t_{i},Z_{i}=z_{i})|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\right]$. Since an entity can have an arbitrary number of neighbors, a GNN is employed to learn an aggregation of the neighbors’ features. Jiang and Sun (2022) shows that if the distribution $p(T_{i}|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}})$, i.e., the classic propensity score, or $p(Z_{i}|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},t_{i})$, i.e., the peer exposure score, is biased (or when both are biased), the estimates for the treatment effects will also be biased. Given that treatments are typically not allocated randomly in observational data due to confounding, controlling for these distributions is necessary.
To address the challenge of confounding bias in the presence of interference, representation balancing techniques can be applied (Ma and Tresp, 2021; Jiang and Sun, 2022; Shalit et al., 2017). This leads to more accurate estimates of the ITTE. In the experiments presented in Section 5, we adopt NetEst (Jiang and Sun, 2022), a GNN-based causal estimator that employs an adversarial approach to balance representations. Nevertheless, alternative relational causal estimators could also be employed.
@@ 4.2 Allocating treatments optimally
Assuming we can compute unbiased ITTE estimates for all entities as $\hat{\omega}_{i}(t_{i},z_{i})=\mathcal{M}(\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},t_{i},z_{i})$- $\mathcal{M}(\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},0,0)$, we can plug these estimates into the objective (Equation 5). As already explained, the strategy of ranking entities according to estimated ITTE no longer works. To address this, we rely on heuristics developed in the IM literature and plug data-driven estimates of TTE into these algorithms. Generally, these heuristics maximize TTE by iteratively improving a (set of) solution(s). Some examples include a greedy algorithm (Kempe et al., 2003), a genetic algorithm (Bucur and Iacca, 2016), and simulated annealing (Jiang et al., 2011). OTAPI can use any of these algorithms to optimize treatment allocation.
@@ 5.1.1 Data
Synthetic data.  For each synthetic dataset, 3 separate networks (train, validation, and test) are generated according to the Barabási-Albert random network model (Barabási and Albert, 1999). Each network consists of 5,000 nodes. The hyperparameter which determines the number of edges attached from each new node is set to $m=2$. The data-generating process (DGP) is inspired by Jiang and Sun (2022).
For each node, a feature vector of $10$ features is randomly sampled according to a standard normal distribution: ${x}_{i}^{j}\sim\mathcal{N}(0,1),j=1,\dots,10$. We first generate the parameters that will be used later on in the DGP:
@@
where Unif is the uniform distribution. These parameters are used to quantify the effect of $\mathbf{X_{i}}$ on ${T_{i}}$, $\mathbf{X_{i}}$ on $Y_{i}$, and ${T_{i}}$ on $Y_{i}$, respectively. The treatment $t_{i}$ is generated as follows:
$t_{i}\sim\text{Bernoulli}(p_{t_{i}});\quad p_{t_{i}}=\sigma(\beta_{XT}\cdot p_{i}+\beta_{{\mathcal{N}}T}\cdot p_{\mathcal{N}_{i}}),$
@@
with $p_{i}=\mathbf{w}^{XT}\cdot\mathbf{x}_{i}$ and $p_{\mathcal{N}_{i}}=\frac{\sum_{j\in\mathcal{N}_{i}}p_{j}}{|\mathcal{N}_{i}|}$. Here, $\sigma$ is the sigmoid function. Note that confounding is introduced by generating propensity scores $p_{i}$ and $p_{\mathcal{N}_{i}}$ based on $\mathbf{x}_{i}$ and $\mathbf{x}_{\mathcal{N}_{i}}$, respectively. Now, we describe how the outcome $y_{i}$ is generated:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}y^{prob}_{i}=\sigma(\beta_{0}+h_{i}\cdot\beta_{{individual}}\cdot t_{i}+h_{i}\cdot\beta_{{spillover}}\cdot z_{i}+\beta_{XY}\cdot u_{i}+\beta_{\mathcal{N}Y}\cdot u_{\mathcal{N}_{i}}+\beta_{\epsilon}\cdot\epsilon);\quad\epsilon\sim\mathcal{N}(0,1),}$
@@
where $h_{i}=\sigma(\mathbf{w}^{TY}\cdot\mathbf{x}_{i})+b_{TY}$, $u_{i}=\sigma(\mathbf{w}^{XY}\cdot\mathbf{x}_{i})$, and $u_{\mathcal{N}_{i}}=\frac{\sum_{j\in\mathcal{N}_{i}}u_{j}}{|\mathcal{N}_{i}|}$. Recall that $z_{i}$ is the ratio of $i$’s neighbors that received treatment $z_{i}=\frac{\sum_{j\in\mathcal{N}_{i}}t_{j}}{|\mathcal{N}_{i}|}$. Finally, $y_{i}$ is sampled using a Bernoulli distribution $y_{i}\sim\text{Bernoulli}(y_{i}^{prob})$. $h_{i}$ is used to make the treatment effect heterogeneous, i.e., dependent on the features. Note that the individual effect is 100% correlated with the spillover effect. This choice is motivated by the fact that if someone is more/less susceptible to treatment, they will most likely also be more/less susceptible to the treatment of neighbors. The following setting for the parameters was used for all experiments: $\beta_{0}=-3$, $\beta_{{individual}}=1$, $\beta_{XT}=1$, $\beta_{\mathcal{N}T}=0.5$, $\beta_{XY}=0.7$, $\beta_{\mathcal{N}Y}=0.2$, $\beta_{\epsilon}=0.05$, and $b_{TY}=4$. The intuition behind these parameters is that the individual features and treatments will always be more impactful than those of neighbors. $\beta_{spillover}$ is varied between 0 and 0.7 across experiments.
Semi-synthetic data.Following Jiang and Sun (2022) and Guo et al. (2020), we use the BlogCatalog (BC) and Flickr datasets to construct semi-synthetic data: the network structures are derived from the real-world data, while both the treatments and outcomes are simulated. Additionally, we use the Enron email interaction dataset (Rossi and Ahmed, 2015) to generate another semi-synthetic dataset.
BC is a social network platform where people can share their blogs. In the BC dataset, the bloggers are the nodes, and the edges represent whether two bloggers are friends on the platform. The node features are bag-of-words vectors derived from the keywords in the bloggers’ textual descriptions. Hence, the features are sparse and high-dimensional. Therefore, following Guo et al. (2020), we use Latent Dirichlet Allocation (LDA) (Blei et al., 2003) to reduce the dimensionality, and construct more informative features. Product sales are simulated as the outcome variable, influenced by a targeted marketing campaign that represents the treatment. Based on their features, some bloggers may be more susceptible to the campaign than others. Additionally, if bloggers start posting or talking about the product of interest, word-of-mouth effects may amplify its visibility and drive product sales.
Flickr is a social media platform where users can share images and videos. In the corresponding dataset, each node represents a user, and edges denote friendship links between users. Node features capture the tags associated with each user’s content interests. As with the BC dataset, these features are represented as high-dimensional, sparse vectors. Hence, the same approach (LDA) is adopted to reduce the dimensionality. The downstream task is defined analogously to the BC case: the treatment is a marketing campaign intended to promote a product, and the outcome is product sales.
In the Enron e-mail interaction dataset, each node represents an employee, and edges correspond to email communications between them. Social ties among employees are a key factor in turnover (Soltis et al., 2013), which can be socially contagious—when one employee leaves, the likelihood that their close colleagues will also leave may increase (Porter and Rigby, 2021). This highlights the importance of the social network in understanding turnover dynamics. Identifying employees with high network impact is therefore crucial, as retaining them—through, for example, salary increases, additional benefits, or professional development opportunities—can help prevent cascading departures and maintain organizational stability. Since no node features are provided, we generate them using node2vec (Grover and Leskovec, 2016), which produces feature embeddings based on the network structure. In this dataset, the treatment corresponds to offering a pay raise, and the outcome is whether the employee eventually leaves the organization (i.e., turnover).
We follow the dataset construction procedure described in Jiang and Sun (2022). To evaluate the out-of-sample performance of OTAPI, each network is partitioned into training, validation, and test sets using the METIS graph partitioning algorithm (Karypis and Kumar, 1998). For the Flickr and BC datasets, we use the same data splits as in Jiang and Sun (2022). The feature dimension of these two datasets is reduced to 10. To assess OTAPI’s performance with more node features, the feature dimension for the Enron dataset is set to 100. Summary statistics for all semi-synthetic datasets are presented in Table 1. After feature generation, the same DGP as in the synthetic datasets is used. For the semi-synthetic datasets, we set $\beta_{spillover}=0.3$; all other parameters remain unchanged.
Flickr BC Enron
Train Validation Test Train Validation Test Train Validation Test
Nodes 2,482 2,461 2,358 1,716 1,696 1,784 11,213 11,232 11,232
Edges 46,268 14,419 23,529 17,937 25,408 14,702 33,228 55,046 63,027
Features 12,047 12,047 12,047 8,189 8,189 8,189 100 100 100
Table 1: Summary statistics for the Flickr, BC, and Enron datasets.
@@
Figure 3: NetEst architecture.